1. What volume of a 6.0 M NaOH solution is needed to produce 3.0 L of a 0.1 M solution?
1
Expert's answer
2021-07-20T14:17:02-0400

Needed n(NaOH) "= 0.1 \times 3.0 = 0.3 \;mol"

Proportion:

6.0 mol – 1000 mL

0.3 mol – x

"x = \frac{0.3 \times 1000}{6.0}=50 \;mL"

Answer: 50 mL